# Identifiability of VAR(1) model in a stationary setting

### v2

@@ 5.3 Summary and illustration
In this section, we first provide a summary of all the identifiability results. In particular, conditions for two stationary VAR(1) models $\mathbf{M}_{1}$ and $\mathbf{M}_{2}$ to satisfy the identifiability criteria or not are listed below:
Table 1: A summary of identifiability results
@@
$\mathbf{M}_{1}$ and $\mathbf{M}_{2}$ are identifiable $G_{1}$ and $G_{2}$ do not contain multi-edges, ${\operatorname{dim}}\left(\mathbf{M}_{1}\right)={\operatorname{dim}}\left(\mathbf{M}_{2}\right)$, and $\mathfrak{MC}_{1}\neq\mathfrak{MC}_{1}$ Theorem 3 $G_{1}$
@@
$\mathbf{M}_{1}$ and $\mathbf{M}_{2}$ do not satisfy the identifiability criteria in this paper ${\operatorname{dim}}\left(\mathbf{M}_{1}\right)={\operatorname{dim}}\left(\mathbf{M}_{2}\right)$ and $\mathfrak{MC}_{1}=\mathfrak{MC}_{2}$ $G_{1}$
@@
Here, $\mathfrak{MC}_{i}$ is the set of maximal classes of the model $\mathbf{M}_{i}$. Note that although we present cases where the identifiability criteria used in this paper are not satisfied, it does not necessarily mean that the models are not identifiable. These are open cases left for future work.
@@
$G_{1}$$\left\{\{1,2\},\{1,3\},\{1,4\}\right\}$ $7$
$G_{2}$$\left\{\{1,3\},\{2,3,4\}\right\}$ $8$
$G_{3}$$\left\{\{1,2\},\{2,3,4\}\right\}$ $8$
@@ 6.1 Bipartite graphs with prior classification
If we already know that species $1-5$ are resources, and species $6-10$ are consumers, then each maximal class contains exactly one resource and all the consumers it feeds, thus the whole graph is identifiable.
@@ 6.3 Resilience of the network
In this case, $A$ is more resilient than $B$, because the only two maximal classes of $A$ are disjoint, and hence nodes $1-3$ do not affect nodes $4-6$ at all. In particular, if we take node $3$ away from the system, then in $A$, only nodes $1$ and $2$ will be affected, while in $B$, all the other nodes will be affected.
@@ Acknowledgments
Appendix A Proof of the fact that the parametrization map is rational
@@
Appendix B Proof of Lemma 1
@@
If we order the rows and columns of $\overline{\mathbf{J}}$ such that the rows correspond to $\lambda_{11}$,$\lambda_{12}$,…,$\lambda_{1n}$,…,$\lambda_{n1}$,$\lambda_{n2}$,…,$\lambda_{nn}$, and $\omega$ respectively, and the columns correspond to $\sigma_{11}$,$\sigma_{12}$,…,$\sigma_{1n}$,…,$\sigma_{n1}$,$\sigma_{n2}$,…,$\sigma_{nn}$, then we are able to organize $A$ and $B$ into the following matrix forms:
@@
Appendix C Proof of Lemma 2
@@
These equations show that elements in $\overline{\mathbf{J}}_{G}$ are exactly the corresponding elements in $\overline{\mathbf{J}}$ setting $\lambda_{st}$ to 00 for all $s,t\in\left[n\right]$ s.t. $\left(s,t\right)\notin\mathfrak{E}_{G}$. ∎
Appendix D Proof of Lemma 3
@@
Appendix E Proof of Lemma 4
@@
Consider any node in the graph, if it has in-degree 00, then it is a source node, and belongs to the maximal class with itself being the source node. If it has a positive in-degree, then it belongs to the same maximal class as its parent node. ∎
@@
Appendix F Proof of Lemma 5
@@
$M^{S_{1}}$ :=$:=\left\{\Lambda\in M_{n}\left(\mathbb{R}\right)\mid\forall i,j\in V,\sum_{k=0}^{+\infty}\sum_{a=1}^{n}\lambda_{ai}^{\left[k\right]}\lambda_{aj}^{\left[k\right]}=0\Leftrightarrow\forall k\in\mathbb{N},a\in\left[n\right],\lambda_{ai}^{\left[k\right]}\lambda_{aj}^{\left[k\right]}=0\right\},$
$M^{S_{2}}$ :=$:=\left\{\Lambda\in M_{n}\left(\mathbb{R}\right)\mid\forall i,j\in V,\sum_{s=1}^{n}\sum_{k=0}^{+\infty}\sum_{l=1}^{n}\lambda_{li}^{\left[k\right]}\lambda_{ls}^{\left[k\right]}\lambda_{sj}=0\Leftrightarrow\forall k\in\mathbb{N},s,l\in\left[n\right],\lambda_{li}^{\left[k\right]}\lambda_{ls}^{\left[k\right]}\lambda_{sj}=0\right\}.$
@@
because the set of roots of a non-zero real analytic function has Lebesgue measure zero ([5, Section 3.1]). This implies that
@@
Appendix G Proof of Lemma 7
@@
Appendix H Proof of Lemma 9
@@
where $M_{G}$ is a diagonal matrix whose elements on the diagonal are $\left(\Sigma\Lambda\right)_{jj}$, s.t. $\exists i\in\left[n\right],\left(i,j\right)\in\mathfrak{E}_{G}^{\prime}$ and $-2$. Hence $M_{G}$ is invertible, and
@@
Appendix I Proof of Case 2 in Theorem 5
@@
Consider a $E_{G}\times E_{G}$ submatrix of $\psi_{G}(B)$, where the rows correspond to $\lambda_{ij}$ s.t. $i,j\in[n]$ and $(i,j)\in\mathfrak{E}_{G}$, and the columns correspond to $\sigma_{ab}$ s.t. $a,b\in[n]$, $a\leq b$, and $(a,b)$ or $(b,a)\in\mathfrak{E}_{G}$. In fact, this submatrix contains all distinct non-zero columns of $\psi_{G}(B)$. Therefore, if this submatrix is full rank, then $\mbox{rank}\left(\psi_{G}\left(B\right)\right)=n_{c}^{\prime}=E_{G}$.

### v3

@@ 5.3 Summary and illustration
In this section, we first provide a summary of all the identifiability results (Table 1), in which conditions for two stationary VAR(1) models $\mathbf{M}_{1}$ and $\mathbf{M}_{2}$ to satisfy the identifiability criteria or not and corresponding examples are listed. In the table, $\mathfrak{MC}_{i}$ is the set of maximal classes of the model $\mathbf{M}_{i}$.
Table 1: A summary of identifiability results with examples
@@
$\mathbf{M}_{1}$ and $\mathbf{M}_{2}$ are identifiable $G_{1}$ and $G_{2}$ do not contain multi-edges, ${\operatorname{dim}}\left(\mathbf{M}_{1}\right)={\operatorname{dim}}\left(\mathbf{M}_{2}\right)$, and $\mathfrak{MC}_{1}\neq\mathfrak{MC}_{2}$ Theorem 3 $G_{1}$
@@
$\mathbf{M}_{1}$ and $\mathbf{M}_{2}$ do not satisfy the identifiability criteria in this paper ${\operatorname{dim}}\left(\mathbf{M}_{1}\right)={\operatorname{dim}}\left(\mathbf{M}_{2}\right)$ and $\mathfrak{MC}_{1}=\mathfrak{MC}_{2}$ - $G_{1}$
@@
Note that although we present cases where the identifiability criteria are not satisfied, this does not necessarily mean the models are unidentifiable. These graphs fall into two categories: (i) those containing multi-edges, the dimensions of which cannot be determined using the results in this paper (final row, Table 1); and (ii) those that cannot be distinguished by model dimension or maximal classes (second-to-last row, Table 1). The latter case could suggest a certain degree of homogeneity.
@@
$G_{1}$ $\left\{\{1,2\},\{1,3\},\{1,4\}\right\}$ 77
$G_{2}$ $\left\{\{1,3\},\{2,3,4\}\right\}$ 88
$G_{3}$ $\left\{\{1,2\},\{2,3,4\}\right\}$ 88
@@ 6.1 Bipartite graphs with prior classification
If we already know that species 151-5 are resources, and species 6106-10 are consumers, then each maximal class contains exactly one resource and all the consumers it feeds, thus the whole graph is identifiable.
@@ 6.3 Resilience of the network
In this case, $A$ is more resilient than $B$, because the only two maximal classes of $A$ are disjoint, and hence nodes 131-3 do not affect nodes 464-6 at all. In particular, if we take node $3$ away from the system, then in $A$, only nodes $1$ and $2$ will be affected, while in $B$, all the other nodes will be affected.
@@ Acknowledgments
Appendix A Algebraic dimension
This appendix introduces formally the algebraic dimension with respect to Zariski topology in the framework of algebraic geometry ([9]), and explains why the dimension can be defined as the rank of the Jacobian matrix at a generic point.
First, we introduce the notion of variety, on which the algebraic dimension is defined.
Definition 11.
Let $\mathbb{R}\left[X\right]$ be the set of polynomials of $n$ variables over $\mathbb{R}$, where $X\in\mathbb{R}^{n}$, and $S\subseteq\mathbb{R}\left[X\right]$ be a subset. The variety defined by $S$ is the set:
$\mathcal{V}(S):=\left\{X\in\mathbb{R}^{n}\mid f(X)=0,\forall f\in S\right\}.$
The variety $V(S)$ is also called the zero set of $S$. A variety of algebraic geometry is not always a manifold in differential geometry because it could contain singularities (e.g. sharp points). Conversely, varieties without singular points fall into the domain of differential geometry as a smooth manifold.
Now, we define the algebraic dimension of a variety.
Definition 12.
Let $V$ be a variety. Define the dimension of $V$ (denoted as ${\operatorname{dim}}(V)$) to be the supremum of all integers $n$ such that there exists a chain $V_{0}\subsetneq V_{1}\subsetneq\cdots\subsetneq V_{n}$ of distinct subvarieties of $V$.
This definition corresponds to Zariski topology, in which the closed sets are defined as the zero sets of any set of polynomials.
Example 14.
The dimension of $\mathbb{R}^{1}$ is $1$, ([9, Chapter 1]). Indeed, the only subvarieties of $\mathbb{R}^{1}$ are the whole space and single points.
The dimension of non-variety sets is simply the dimension of its algebraic closure, i.e. the smallest variety containing the set. For example, $\mathbb{R^{+}}$ is not a variety. Its dimension is $1$ because its algebraic closure is $\mathbb{R}^{1}$.
Our model $\mathbf{M}_{G}$ is also not a variety, the dimension would refer to the dimension of its algebraic closure. This dimension is calculable by Gröbner bases ([18, Section 3.4]), yet computationally demanding. Instead, we use an equivalent definition for the dimension - the rank of the Jacobian matrix of the parametrization map at a generic point. This is true because the parametrization map is rational (proof in Appendix B) ([18, Chapter 16, Th.16.1.7]).
Appendix B Proof of the fact that the parametrization map is rational
@@
Appendix C Proof of Lemma 1
@@
If we order the rows and columns of $\overline{\mathbf{J}}$ such that the rows correspond to $\lambda_{11}$$\lambda_{12}$$\lambda_{1n}$$\lambda_{n1}$$\lambda_{n2}$$\lambda_{nn}$, and $\omega$ respectively, and the columns correspond to $\sigma_{11}$$\sigma_{12}$$\sigma_{1n}$$\sigma_{n1}$$\sigma_{n2}$$\sigma_{nn}$, then we are able to organize AA and $B$ into the following matrix forms:
@@
Appendix D Proof of Lemma 2
@@
These equations show that elements in $\overline{\mathbf{J}}_{G}$ are exactly the corresponding elements in $\overline{\mathbf{J}}$ setting $\lambda_{st}$ to 0 for all $s,t\in\left[n\right]$ s.t. $\left(s,t\right)\notin\mathfrak{E}_{G}$. ∎
Appendix E Proof of Lemma 3
@@
Appendix F Proof of Lemma 4
@@
Consider any node in the graph, if it has in-degree 0, then it is a source node, and belongs to the maximal class with itself being the source node. If it has a positive in-degree, then it belongs to the same maximal class as its parent node. ∎
@@
Appendix G Proof of Lemma 5
@@
$M^{S_{1}}$ $:=\left\{\Lambda\in M_{n}\left(\mathbb{R}\right)\mid\forall i,j\in V,\sum_{k=0}^{+\infty}\sum_{a=1}^{n}\lambda_{ai}^{\left[k\right]}\lambda_{aj}^{\left[k\right]}=0\Leftrightarrow\forall k\in\mathbb{N},a\in\left[n\right],\lambda_{ai}^{\left[k\right]}\lambda_{aj}^{\left[k\right]}=0\right\},$
$M^{S_{2}}$ $:=\left\{\Lambda\in M_{n}\left(\mathbb{R}\right)\mid\forall i,j\in V,\sum_{s=1}^{n}\sum_{k=0}^{+\infty}\sum_{l=1}^{n}\lambda_{li}^{\left[k\right]}\lambda_{ls}^{\left[k\right]}\lambda_{sj}=0\Leftrightarrow\forall k\in\mathbb{N},s,l\in\left[n\right],\lambda_{li}^{\left[k\right]}\lambda_{ls}^{\left[k\right]}\lambda_{sj}=0\right\}.$
@@
because the set of roots of a non-zero real analytic function has Lebesgue measure zero ([6, Section 3.1]). This implies that
@@
Appendix H Proof of Lemma 7
@@
Appendix I Proof of Lemma 9
@@
where $M_{G}$ is a diagonal matrix whose elements on the diagonal are $\left(\Sigma\Lambda\right)_{jj}$, s.t. $\exists i\in\left[n\right],\left(i,j\right)\in\mathfrak{E}_{G}^{\prime}$ and 2-2. Hence $M_{G}$ is invertible, and
@@
Appendix J Proof of Case 2 in Theorem 5
@@
Consider a $E_{G}\times E_{G}$ submatrix of $\psi_{G}(B)$, where the rows correspond to $\lambda_{ij}$ s.t. $i,j\in[n]$ and $(i,j)\in\mathfrak{E}_{G}$, and the columns correspond to $\sigma_{ab}$ s.t. $\in[n], aba\leq b$, and $(a,b)$ or $(b,a)\in\mathfrak{E}_{G}$. In fact, this submatrix contains all distinct non-zero columns of $\psi_{G}(B)$. Therefore, if this submatrix is full rank, then $\mbox{rank}\left(\psi_{G}\left(B\right)\right)=n_{c}^{\prime}=E_{G}$.